DocDB 8816 April 2013 Muon Operations Summary Water Systems From James Wilhelmi of RPI (DocDB 8943): Water system operations - April 2013 For meeting on May 3, 2013 EH4: Continues to operate within specification, for the most part. Frequent instrument alarms associated with the low RO pH have been a notable issue. EH1/2/3 April 12 SS vacuum tanks arrived onsite. EH1 was installed, EH2 and EH3 are not yet installed. Sporadic readings have been taken with the auxiliary resistivity sensors. Operating with 1 and 0.2 micron filters. All the dissolved oxygen sensors are need to be calibrated and are possible faulty, their readings do not seem trustworthy. Random system stops continue to occur, for reasons unknown. EH1 April 12 - New resistivity sensor ports where installed before the vacuum tank and after the heat exchanger. Pool water DCS sensor now currently resides in the new port before the vacuum tank. EH1 April 12 - Level sensors installed in March were fully integrated with the arrival of a new PLC unit. Muon PMTs As reported previously, we would like to replace some EH3 PMTs, but it is unlikely that there will be any opportunities for this. There is a visible reduction in pool veto efficiency at the location corresponding to two dead, face-to-face PMTs. See DocDB 8858. Dead PMTs as of March, 2013: EH1: DVIF42 DVIB32 DVIB41 EH2: LCIG72 LVIA44 LVOEF2 EH3: FCIA61 FVOE32 FVIE41 FVOA43 FVIA43 FVOC43 FVOG31 FVIG41 All of the failures are Macros. Muon Calibration From Jo Ellen of VT: Status: o "Full" water pool calibration for three halls run this month. o http://web.dyb.ihep.ac.cn:8099/Calibration/96 o http://web.dyb.ihep.ac.cn:8099/Calibration/98 o http://web.dyb.ihep.ac.cn:8099/Calibration/100 o http://web.dyb.ihep.ac.cn:8099/Calibration/102 o http://web.dyb.ihep.ac.cn:8099/Calibration/104 o Known issue: On Apr. 26, during EH3 WP Calibration, the WPCalib software has been reset automatically or by accident twice. The possible reason is either a lost network connection or a PC memory allocation problem. The details are described in elog entries 104. Investigation continues. o Known issue: The EH3 OWS PMT-like signal (reference signal) can't be read by FEE(Board15, Channel7(counting from 1)). Need to check the cable connection and FEE B15C7 channel readout. Actions and Plan: o For the EH3 OWS PMT-like signal channel, need to check the cable connection and FEE B15C7 channel readout. If FEE B15C7 can't read the signal, switch to B12C15, which is a dead OWS PMT channel, and suggested by Bob. If we get permission, we will send a detailed procedure to coordinator and shifter. RPC Performance No issues. RPC Gas System No issues. RPC HV From Jo Ellen of VT: Status: o Running smoothly Action and plan: o None Overall Muon Performance Performance continues unchanged.